OpenStudy (anonymous):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 40 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. f(x) = - 2x + 40 f(x) = 2x + 40 f(x) = - 2x + 50 f(x) = 2x + 50

OpenStudy (anonymous):

translate as find the equation of the line through \((5,40)\) and \((10,30)\)

OpenStudy (anonymous):

slope is \(\frac{30-40}{10-5}=\frac{-10}{5}=-2\)

OpenStudy (yanasidlinskiy):

You would first have to calculate the slope.

OpenStudy (anonymous):

@YanaSidlinskiy and @satellite73 are correct. but i have this my own way plug (10,30) in each equation, only 1 equation satisfies the condition. this way will not work always... but hopefully it worked here :)

OpenStudy (yanasidlinskiy):

You can already eliminate 2 choices or 3. You have your slope as -2. You can rule out the positives and stick with the negative. Now you're down to 50/50:D

OpenStudy (anonymous):

thanks guiseeeee !

OpenStudy (anonymous):

welcome :D

OpenStudy (anonymous):

Was it A or C

OpenStudy (anonymous):

@Jasmineskyee c. f(x) = - 4x + 80